OpenStudy (anonymous):

4. A train is travelling at a constant speed of 15 m/s. It starts to accelerate with an acceleration of 1 m/s2 for 3 seconds. What is the distance travelled by the train during the 3 seconds interval when it is accelerating?

OpenStudy (dumbcow):

do you know calculus

OpenStudy (anonymous):

No, cause we are learning physics.

OpenStudy (anonymous):

>.>

OpenStudy (dumbcow):

then you have to approximate with time intervals For 1st second , the speed will increase from 15 to 16m/s 2nd second, increase from 16 to 17 3rd second, increase from 17 to 18

OpenStudy (dumbcow):

im not exactly sure on what method your teacher or class wants this to be done though

OpenStudy (anonymous):

Alright. Then may i know what is the next step?

OpenStudy (anonymous):

is tgis constant accelaration for 3 seconds?

OpenStudy (anonymous):

if so we can use newtons equations for constant acceleration

OpenStudy (dumbcow):

ok pick a midpoint for each interval which we will use to approximate for the constant speed during a 1-sec interval. 1st second : 15.5 2nd second: 16.5 3rd second: 17.5 distance = rate*time, time is 1 for each case Add those up to get distance

OpenStudy (anonymous):

s = ut + 1/2 at^2 = =15*3 + 1.2 * 1 * 3^2 = 45 + 5.5 = 49.5 meters

OpenStudy (anonymous):

agrees with dumbcow's answer

OpenStudy (dumbcow):

well yeah i could've used that too but it involved calculus to derive it

OpenStudy (anonymous):

May i know why i get the answer if i add up the midpoint ?

OpenStudy (anonymous):

And may i know what formula is this = ut + 1/2 at^2?

OpenStudy (dumbcow):

because say for 1st second, the speed is constantly changing right, however the acceleration is constant so if you take the speed right in the middle and apply to entire interval it captures all the changes. because early in the interval the speed is higher than actual but later on the speed will be lower than actual, they kind of cancel each other out. hope that made some sense

OpenStudy (anonymous):

Alright. Thank you !

OpenStudy (anonymous):

these equations were developed using calculus as dumbcow said they apply to motion at constant accekleration only there are 4 main ones s = (u + v )t / 2 v = u + at s = ut + (1/2) at^2 v^2 = u^2 + 2as where s = distance, u = initial velocity, v = final velocity, a = accelerartion, t=time

OpenStudy (anonymous):

So actually , if the qn is asking for constant acceleration i can apply these four fomulas?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they are very useful for falling objects where a = acceleration due to gravity they are usually introduced in physics before calculus is taught ( in UK that is)